xx TestNorms

TestNorms

Roland Haas <rhaas@illinois.edu>

December 28 2020

Abstract

1 Introduction

2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Examples 4.6 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References 6 Parameters

11 refinedxradius10 Scope: private01  REAL

11 Description: size of the refined region at the center

11 Range 10              01  Default: 0.25
11 (0:*  10 any positive size

11

7 Interfaces

General Implements: testnorms

Inherits: driver carpetxregrid

Grid Variables

7.0.1 PRIVATE GROUPS

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11  gf000  10  gf000  01  centering 11 centering={0 0 0}

11                    10                    01  compact 11 0

11                    10                    01 description 11 Test grid function
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  gf001  10  gf001  01  centering 11 centering={0 0 1}

11                    10                    01  compact 11 0

11                    10                    01 description 11 Test grid function
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  gf010  10  gf010  01  centering 11 centering={0 1 0}

11                    10                    01  compact 11 0

11                    10                    01 description 11 Test grid function
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  gf011  10  gf011  01  centering 11 centering={0 1 1}

11                    10                    01  compact 11 0

11                    10                    01 description 11 Test grid function
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  gf100  10  gf100  01  centering 11 centering={1 0 0}

11                    10                    01  compact 11 0

11                    10                    01 description 11 Test grid function
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  gf101  10  gf101  01  centering 11 centering={1 0 1}

11                    10                    01  compact 11 0

11                    10                    01 description 11 Test grid function
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL

11

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11  gf110  10  gf110  01  centering 11 centering={1 1 0}

11                    10                    01  compact 11 0

11                    10                    01 description 11 Test grid function
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  gf111  10  gf111  01  centering 11 centering={1 1 1}

11                    10                    01  compact 11 0

11                    10                    01 description 11 Test grid function
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL

Uses header: loop.hxx

8 Schedule This section lists all the variables which are assigned storage by thorn CarpetX/TestNorms. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage x

11 Always: 11 x
11 gf000  11 x
11 gf001  11 x
11 gf010  11 x
11 gf011  11 x
11 gf100  11 x
11 gf101  11 x
11 gf110  11 x
11 gf111  11 x

11 x  11 x

11

Scheduled Functions

CCTKxPOSTINITIAL x testnormsxseterror x set up test grid x

11 x 11 Language: 11 c
11 x 11 Type:  11 function
11 x 11 Writes: 11 carpetxregrid::regridxerror(interior)

11

CCTKxINITIAL x testnormsxset x set up test data

11 x 11 Language: 11 c
11 x 11 Type:  11 function
11 x 11 Writes: 11 gf000(all)
11 x 11 x  11 gf001(all)
11 x 11 x  11 gf010(all)
11 x 11 x  11 gf011(all)
11 x 11 x  11 gf100(all)
11 x 11 x  11 gf101(all)
11 x 11 x  11 gf110(all)
11 x 11 x  11 gf111(all)